OpenStudy (anonymous):

5x²+5x-12 / (x-1) = 0

OpenStudy (lgbasallote):

solve for x?

OpenStudy (anonymous):

yup do you do LCD ?

OpenStudy (lgbasallote):

no..you just cross multiply

OpenStudy (lgbasallote):

\[5x^2+5x - 12 = 0(x-1)\] \[5x^2 + 5x - 12 = 0\] now factor out and solve for x

OpenStudy (anonymous):

oops typo 2x^2+5x-12=0

OpenStudy (lgbasallote):

then it's \[2x^2 + 5x - 12\] factor out and solve for x

OpenStudy (anonymous):

(2x-x) (x-4)

OpenStudy (anonymous):

(2x-3)*

OpenStudy (lgbasallote):

\[(2x-3)(x-4) = 2x^2 -11x + 12\]

OpenStudy (lgbasallote):

you made a wrong sign somewehre

OpenStudy (anonymous):

yeah (2x+3) (x-4)

OpenStudy (anonymous):

thankss !

OpenStudy (lgbasallote):

welcome :D